OpenStudy (anonymous):

Using complete sentences, compare and contrast the ways to factor the "sum of two squares" and the "difference of two squares." Provide an explanation and example of each, including any similarities and differences.

OpenStudy (anonymous):

Hint, the difference vs. sum are of the form: \[ a^2-b^2=(a+b)(a-b)\\ a^2+b^2=(a+bi)(a-bi) \]